# Nonparametric bounds for vaccine effects in randomized trials

### v1

@@ I.1 Proof that the LP-based bounds in Proposition 5 are sharp
Appendix N Additional results for the numerical example
@@
Figure N.1: The relationship between the magnitude of $\gamma_{B^{1}}$ and the width of the bounds for $\mathrm{VE}(0)$, $\mathrm{VE}(1)$, and $\mathrm{VE}_{T}$ under the setting of Figure 3(d). Assumption 8(i) is preserved when $\gamma_{B^{1}}>0$ and increasingly violated as $\gamma_{B^{1}}$ becomes negative, with $\gamma_{B^{0}}>0$ throughout. The parameters $\gamma_{B^{0}}$ and $\gamma_{B^{1}}$ control whether this assumption holds: when both were positive, the assumption is preserved, whereas more negative values of $\gamma_{B^{1}}$ correspond to stronger violations. Circles represent the true values. LP: bounds constructed via Proposition 6 and the results in Appendix E; Monotonicity: bounds constructed via Proposition 7. Triangles at the left boundary indicate bounds that extended below 0% but were truncated for display. No bounds are displayed when the estimated lower bound exceeds the upper bound, reflecting violation of the assumed monotonicity assumption.
@@
Figure N.2: The relationship between the magnitude of $\delta_{U}$ and the width of the bounds for $\mathrm{VE}(0)$, $\mathrm{VE}(1)$, and $\mathrm{VE}_{T}$ under the setting of Figure 3(c) ($\beta_{S}=0$). The case $\delta_{U}=0$ corresponded to the situation where our assumed DAG matched the true data-generating DAG. As $\delta_{U}$ increased, however, the true data moved further away from this assumed structure, since the effect of $U$ on $S$ became stronger (Figures 3(a) and 3(c)).Circles represent the true values. LP-based bounds constructed via Proposition 4 and the results in Appendix E. Triangles at the left boundary indicate bounds that extended below 0% but were truncated for display.
@@
Figure N.3: The relationship between the magnitude of $\delta_{U}$ and the width of the bounds for $\mathrm{VE}(0)$, $\mathrm{VE}(1)$, and $\mathrm{VE}_{T}$ under the setting of Figure 3(d). Here $\beta_{S}=log(2)$, so the data were generated with both $U$ and $S$ affecting $Y$, while the assumed DAG omitted these effects (Figure 3(a)). Circles represent the true values. LP-based bounds constructed via Proposition 4 and the results in Appendix E. Triangles at the left boundary indicate bounds that extended below 0% but were truncated for display. No bounds are displayed when the estimated lower bound exceeds the upper bound, reflecting violation of the assumed structure.
N.1 Bounds and corresponding bootstrap CIs
Table N.1: Empirical coverage of 95% percentile-bootstrap CIs for LP-based bounds
@@
Table N.2: Empirical coverage of 95% percentile-bootstrap CIs for monotonicity-based bounds
@@
Appendix O Additional results for the real data example
O.1 The full probabilities used to generate $B$ in the real data example
@@
O.2 Additional results for the real data example
Figure O.1: Estimated LP-based and monotonicity-based bounds under Figure 3(a). Each horizontal line shows the identified interval for $VE(0)$, $VE(1)$, and $VE_{T}$, stratified by method (LP vs. Monotonicity) and whether Assumption 8 (i) was imposed (solid = no, dashed = yes). Triangles at the left boundary indicate bounds that extended below 0% but were truncated for display. No bounds are displayed when the estimated lower bound exceeds the upper bound, reflecting violation of the assumed structure.
Table O.1: Estimated VE bounds (%) under Figure 3(d). Cells show Estimate (95% CI). Values below $1000\%$ are reported as -$\infty$.

### v2

@@ Appendix I Proof of Proposition 3: LP-based bounds under Figure 3(a)
Appendix N Narrowing bounds with additional measured covariates
In this section, we use the notation $p^{\,l}_{yb.al}=\Pr(Y=y,B=b\mid A=a,L=l)$, $p^{\,l}_{y.abl}=\Pr(Y=y\mid A=a,B=b,L=l)$, $p^{\,l}_{y.abls}=\Pr(Y=y\mid A=a,B=b,L=l,S=s)$, $\pi^{\,l}_{b.al}=\Pr(B=b\mid A=a,L=l)$, and $\pi^{\,l}_{b.als}=\Pr(B=b\mid A=a,L=l,S=s)$.
N.1 Narrowing bounds with additional measured covariates
In practice, vaccine trials typically collect a set of baseline covariates (other than $S$), denoted by $L$. We begin with the case where certain covariates $L$ can be used to obtain narrower bounds under the conditions described below. When $S$ is unobserved (Figure 2) or cannot be exploited to tighten the bounds (as in Figures 3(b)3(d)), such covariates $L$ may serve this purpose. When conditioning on $S$ leads to narrower bounds (as in Figure 3(a)), replacing stratification by $S$ alone with joint stratification by $(S,L)$ may further narrow them. For illustration, we present only the LP-based bounds for $VE(0)$ in the main text. The remaining LP-based bounds, the monotonicity-based bounds, and the full proofs are deferred to the rest of Appendix N.
From Section 5.1, we know that $S$ can be used to obtain narrower bounds if it satisfies $S\perp_{\mathcal{T}_{VI}}U$ and $Y\perp_{\mathcal{T}_{VI}}S\mid A,B,M,U$. The analogous result holds for $L$, as summarized in the following proposition.
Proposition 14.
Any covariate $L$ that satisfies $L\perp_{\mathcal{T}_{VI}}(A,M,S,U)$ and $Y\perp_{\mathcal{T}_{VI}}L\mid A,S,B,M,U$ can be used to obtain narrower bounds under both the LP-based and the monotonicity-based approaches.
An example of an $L$ that might satisfy the above conditions is participants’ awareness of the randomization ratio. Participants’ awareness of the ratio prior to enrollment could affect their belief about treatment assignment without directly influencing the risk of infection. For instance, in the Novavax COVID-19 vaccine trial, the randomization ratio was 2:1 (vaccine:placebo), and participants were informed of this ratio before enrollment (Dunkle et al., 2022; National Institutes of Health, 2020). In practice, however, not all participants may have been equally attentive to or aware of this information. If only some participants are aware of the randomization ratio (which could be measured by a survey), conditioning on awareness can be used to obtain narrower bounds. An example of an $L$ that fails to meet the conditions above is when $L$ represents the observed part of $U$, that is, a measured common cause of both $B$ and $Y$. In this case, as well as in the other cases illustrated in Figure N.2, $L$ cannot be used to obtain narrower bounds. Further discussion with real data examples appears in Appendix N. In practice, most baseline covariates in vaccine trials are connected to $U$, $Y$, or both, and therefore cannot be used to obtain narrower bounds.
When $S$ is unobserved (Figure 2) or does not lead to narrower bounds (Figures 3(b)3(d)), a covariate $L$ satisfying the conditions of Proposition 14 may be used instead. We can apply the results for $S$ in Section 5 and replace $S$ with $L$ in the LP-based bounds of Proposition 3 and in the monotonicity-based bounds of Proposition 4. For example, under Figure 1(a) the LP-based bounds for $VE(0)$ now take the form
$\max_{l_{1},l_{2}\in\{0,1\}^{2}}\left\{1-\frac{1-p^{l}_{00.1l_{1}}}{p^{l}_{10.0l_{2}}}\right\}\leq VE(0)\leq\min_{l_{1},l_{2}\in\{0,1\}^{2}}\left\{1-\frac{p^{l}_{10.1l_{1}}}{1-p^{l}_{00.0l_{2}}}\right\},$
where $p^{l}_{yb.al}=\Pr(Y=y,B=b\mid A=a,L=l)$.
By contrast, when $S$ can be used to obtain narrower bounds (Figure 3(a)), these results can be further refined when $L$ also satisfies Proposition 14 conditions, by replacing stratification by $S$ alone with joint stratification by $(S,L)$. For example, under Figure 1(b) the LP-based bounds for $VE(0)$ now become
$\max_{s_{1},s_{2},l_{1},l_{2}\in\{0,1\}^{4}}\left\{1-\frac{1-p^{l}_{00.1\,s_{1}l_{1}}}{p^{l}_{10.0\,s_{2}l_{2}}}\right\}\;\leq\;VE(0)\;\leq\;\min_{s_{1},s_{2},l_{1},l_{2}\in\{0,1\}^{4}}\left\{1-\frac{p^{l}_{10.1\,s_{1}l_{1}}}{1-p^{l}_{00.0\,s_{2}l_{2}}}\right\},$
where $p^{l}_{yb.asl}=\Pr(Y=y,B=b\mid A=a,S=s,L=l)$.
N.2 Case 1: when $L$ can be used to obtain narrower bounds
As described in Section N.1, when $L$ satisfies the conditions in Proposition 14, it can be used to narrow the LP-based bounds in Proposition 1 and the monotonicity-based bounds in Proposition 2. Examples of such variables $L$ are presented in Figure N.1. In this section, we will proof the bounds for three scenarios.
  1. 1.
    Case 1a. When $L$ can be used to obtain narrower bounds but $S$ cannot be used to obtain narrower bounds (Figure 1(a)).
  2. 2.
    Case 2b. When both $L$ and $S$ can be used to obtain narrower bounds (Figure 1(b)).
  3. 3.
    Case 3c. When $L$ can be used to obtain narrower bounds and $S$ is unobserved (Figure 1(c)).
Although other DAGs besides the one in Figure 1(a) could represent the first scenario, we focus on this DAG and prove the bounds under Figure 1(a), since it is the most “complicated” in the sense that both paths $S\rightarrow Y$ and $S\leftarrow U$ are present.
AASSBBEEUUYYMMLL
(a) Case 1a. Similar to Figure 3(d) but with $L$ connected only to $B$.
AASSBBEELLUUYYMM
(b) Case 1b. Similar to Figure 3(a) but with $L$ connected only to $B$.
AABBEEUUYYMMLL
(c) Case 1c. Similar to Figure 2 but with $L$ connected only to $B$.
Figure N.1: DAGs describing scenarios with observed covariate $L$ that can be used to obtain narower bounds.
N.2.1 A proof that under Proposition 14, $Y_{i}^{a,m}\perp_{\mathcal{T}_{VI}}(A_{i},L_{i})$ holds
For the LP-based bounds, we need to show that if $L$ satisfies the conditions in Proposition 14, then $Y_{i}^{a,m}\perp_{\mathcal{T}_{VI}}(A_{i},L_{i})$ also holds under the NPSEM-IE model. We prove this independence under Figure 1(a); similar arguments apply to Figures 1(b) and 1(c).
Under Figure 1(a), we consider the corresponding NPSEM-IE:
$A_{i}$ $=f_{A}(\varepsilon_{A_{i}}),$ (39)
$L_{i}$ $=f_{L}(\varepsilon_{L_{i}}),$
$U_{i}$ $=f_{U}(\varepsilon_{U_{i}}),$
$M_{i}$ $=f_{M}(\varepsilon_{M_{i}}),$
$S_{i}$ $=f_{S}(A_{i},U_{i},\varepsilon_{S_{i}}),$
$B_{i}$ $=h_{B}(M_{i},S_{i},L_{i},U_{i},\varepsilon_{B_{i}})=\mathbb{I}_{\{M_{i}=-1\}}f_{B}(S_{i},L_{i},U_{i},\varepsilon_{B_{i}})+\mathbb{I}_{\{M_{i}\neq-1\}}M_{i},$
$Y_{i}$ $=f_{Y}(A_{i},B_{i},S_{i},U_{i},\varepsilon_{Y_{i}}),$
for some functions $f_{A},f_{L},f_{U},f_{M},f_{S},f_{B},$ and $f_{Y}$, and where $\varepsilon_{A_{i}},\varepsilon_{L_{i}},\varepsilon_{U_{i}},\varepsilon_{M_{i}},\varepsilon_{S_{i}},\varepsilon_{B_{i}},$ and $\varepsilon_{Y_{i}}$ are mutually independent error terms.
From the model in Equation (39), for $m\in\{0,1\}$, it follows that
$Pr_{II}\Big(Y_{i}^{a,m}=y\,\Big|\,A_{i},L_{i}\Big)$ (40)
$=Pr_{IV}\Big(Y_{i}^{a,m}=y\,\Big|\,A_{i},L_{i}\Big)$
$=Pr_{IV}\Big(f_{Y}(a,B_{i}^{a,m},S^{a}_{i},U_{i},\varepsilon_{Y_{i}})=y\,\Big|\,A_{i},L_{i}\Big)$
$=Pr_{IV}\Big(f_{Y}(a,m,f_{S}(a,f_{U}(\varepsilon_{U_{i}}),\varepsilon_{S_{i}}),f_{U}(\varepsilon_{U_{i}}),\varepsilon_{Y_{i}})=y\,\Big|\,f_{A}(\varepsilon_{A_{i}}),f_{L}(\varepsilon_{L_{i}})\Big)$ (Step 1)
$=Pr_{IV}\Big(f_{Y}(a,m,f_{S}(a,f_{U}(\varepsilon_{U_{i}}),\varepsilon_{S_{i}}),f_{U}(\varepsilon_{U_{i}}),\varepsilon_{Y_{i}})=y\Big)$ (Step 2)
where Step 1 follows from the fact that under Assumption 1 for $m\in\{0,1\}$, $B_{i}^{a,m}=m$. Step 2 follows since $(\varepsilon_{A_{i}},\varepsilon_{L_{i}})\perp(\varepsilon_{U_{i}},\varepsilon_{S_{i}})$.
N.2.2 Case 1a. LP-based bounds
Proposition 15.
Consider the setting of Figure 1(a). The LP-based bounds are
$\max_{l_{1},l_{2}\in\{0,1\}}\!\left\{1-\frac{1-p^{\,l}_{00.1l_{1}}}{p^{\,l}_{10.0l_{2}}}\right\}\leq VE(0)\leq\min_{l_{1},l_{2}\in\{0,1\}}\!\left\{1-\frac{p^{\,l}_{10.1l_{1}}}{1-p^{\,l}_{00.0l_{2}}}\right\},$
$\max_{l_{1},l_{2}\in\{0,1\}}\!\left\{1-\frac{1-p^{\,l}_{01.1l_{1}}}{p^{\,l}_{11.0l_{2}}}\right\}\leq VE(1)\leq\min_{l_{1},l_{2}\in\{0,1\}}\!\left\{1-\frac{p^{\,l}_{11.1l_{1}}}{1-p^{\,l}_{01.0l_{2}}}\right\},$
$\max_{l_{1},l_{2}\in\{0,1\}}\!\left\{1-\frac{1-p^{\,l}_{01.1l_{1}}}{p^{\,l}_{10.0l_{2}}}\right\}\leq VE_{T}\leq\min_{l_{1},l_{2}\in\{0,1\}}\!\left\{1-\frac{p^{\,l}_{11.1l_{1}}}{1-p^{\,l}_{00.0l_{2}}}\right\}.$
Proof. We begin by outlining the main steps of the proof for deriving the LP-based bounds. Let $r_{y\mid a,l,b}$ denote the potential response of $Y^{a,m}$ given $A=a$, $L=l$, and $B^{a,m=-1}=b$, for all $a,m\in\{0,1\}^{2}$ in the two-arm trial. As in Section 4.1, and for clarity, when conditioning on $A=a$ we write $B^{a,m=-1}$ simply as $B$. Furthermore, let $q_{i.alb}=\Pr(r_{y\mid a,l,b}=i)$ and define
$Q_{a,l,b}(a,m)=\{r_{y\mid a,l,b}:\,Y^{a,m}=1\ \text{given }A=a,L=l,B=b\}.$
Then, similarly to Section 4.1, and using the fact that $Y^{a,m}\perp_{\mathcal{T}_{VI}}(A,L)$, we can express the target estimand $E(Y^{a,m})$ as
$\begin{split}E(Y^{a,m})&=E(Y^{a,m}\mid A=a,L=l)\\ &=\pi^{\,l}_{0.al}E(Y^{a,m}\mid A=a,L=l,B=0)+\pi^{\,l}_{1.al}E(Y^{a,m}\mid A=a,L=l,B=1)\\ &=\pi^{\,l}_{0.al}\sum_{i\in Q_{a,l,0}}q^{\,l}_{i.al0}+\pi^{\,l}_{1.al}\sum_{i\in Q_{a,l,1}}q^{\,l}_{i.al1}.\end{split}$ (41)
The constraints and the remaining steps of the proof are identical to those described in Section 4.1, except that $S$ is replaced by $L$.
N.2.3 Case 1a. Monotonicity-based bounds
Consider the following assumption.
Assumption 15.
($U|S,L$ monotonicity) Consider the setting of Figures 1(a). Suppose that $Y\perp_{\mathcal{T}_{VI}}M|A,B,S,U,L$ and the following conditions hold:
  1. 1.
    $\Pr\left(Y=1|A=a,B=1,S=s,L=l,U=u\right)$ and $\Pr(B=1|A=a,S=s,L=l,U=u)$ are both nondecreasing or both nonincreasing in $u$, for all $a$, $s$ and $l$.
  2. 2.
    $\Pr(B=1|A=a,S=s,L=l)$ is nondecreasing in $s$, for all $a$ and $l$.
  3. 3.
    $E\!\left(Y\mid A=a,B=1,S=s,L=l,U\right)$ is nondecreasing in $s$, for all $a$ and $l$.
Proposition 16.
Consider the setting of Figure 1(a). Under Assumptions 8 and 15, the monotonicity-based bounds are
$\max_{l\in\{0,1\}}\!\left\{1-\frac{p_{1.1}}{p^{\,l}_{1.00l}}\right\}$ $\;\leq\;$ $VE(0)$ $\;\leq\;\min_{l\in\{0,1\}}\!\left\{1-\frac{p^{\,l}_{1.10l}}{p_{1.0}}\right\},$
$\max_{l\in\{0,1\}}\!\left\{1-\frac{p^{\,l}_{1.11l}}{p_{1.0}}\right\}$ $\;\leq\;$ $VE(1)$ $\;\leq\;\min_{l\in\{0,1\}}\!\left\{1-\frac{p_{1.1}}{p^{\,l}_{1.01l}}\right\},$
$\max_{l_{1},l_{2}\in\{0,1\}}\!\left\{1-\frac{p^{\,l}_{1.11l_{1}}}{p^{\,l}_{1.00l_{2}}}\right\}$ $\;\leq\;$ $VE_{T}$  $\;\leq\;1-\frac{p_{1.1}}{p_{1.0}}.$
Proof. In Appendix L.3 we proved that
$E(Y^{a,m=1})=\sum_{s}E_{U}\!\left[E\!\left(Y\mid A=a,B=1,S=s,U\right)\gamma_{s.a}|S=s,A=a\right].$
Since under Assumption 15 we assume that $Y\perp_{\mathcal{T}_{VI}}L\mid A,S,B,M,U$, so we can also write that
$E(Y^{a,m=1})=\sum_{s}E_{U}\!\left[E\!\left(Y\mid A=a,B=1,S=s,L=l,U\right)\gamma_{s.a}|S=s,A=a\right].$
From the other side. Under Figure 1(a), we have that $U\perp L|A,S$, thus $E(Y\mid A=a,B=1,L=l)$ can be written as
$\begin{split}E&(Y\mid A=a,B=1,L=l)\\ &=\sum_{s}E\!\left(Y\mid A=a,B=1,L=l,S=s\right)\gamma_{s.abl}\\ &=\sum_{s,u}E\!\left(Y\mid A=a,B=1,S=s,L=l,U=u\right)\gamma_{s.abl}\Pr(U=u\mid A=a,B=1,S=s,L=l)\\ &=\sum_{s,u}E\!\left(Y\mid A=a,B=1,S=s,L=l,U=u\right)\gamma_{s.abl}\frac{\Pr(B=1\mid A=a,U=u,S=s,L=l)\Pr(U=u\mid A=a,S=s,L=l)}{\Pr(B=1\mid A=a,S=s,l=l)}\\ &=\sum_{s,u}E\!\left(Y\mid A=a,B=1,S=s,L=l,U=u\right)\gamma_{s.abl}\Pr(U=u|A=a,S=s)\frac{\Pr(B=1\mid A=a,U=u,S=s,L=l)}{\Pr(B=1\mid A=a,S=s,L=l)}\\ &=\sum_{s}E_{U}\!\left[E\!\left(Y\mid A=a,B=1,S=s,L=l,U\right)\gamma_{s.abl}\frac{\Pr(B=1\mid A=a,S=s,L=l,U)}{\Pr(B=1\mid A=a,S=s,L=l)}\bigg|{A=a,S=s}\right].\end{split}$ (42)
Now, we prove that $E(Y\mid A=a,B=1,L=l)-E(Y^{a,m=1})\geq 0$. Let $\nu_{sl}(u)=\frac{\Pr(B=1\mid A=a,S=s,L=l,U=u)}{\Pr(B=1\mid A=a,S=s,L=l)}$ and $g_{sl}(u)=E\!\left(Y\mid A=a,B=1,S=s,L=l,U=u\right)$. Under Assumption 15, and following the steps in Appendix L.3, we obtain that $E(Y\mid A=a,B=1,L=l)-E(Y^{a,m=1})$ can be written as
[…]
Which is greater or equal to zero since:
  1. 1.
    We assume that $g_{sl}(U)$ is nondecreasing in $u$ and that $\Pr(B=1\mid A=a,S=s,L=l,U=u)$ is nondecreasing in $u$ and thus $\nu_{sl}(u)$ is nondecreasing in $u$. From Lemma 1 it follows that $\operatorname{Cov}_{U}\!\left(g_{s}(U),\,\nu_{s}(U)\right|A=a,S=s)\geq 0$.
  2. 2.
    We assume that $E_{U}(g_{sl}(U)|A=a,S=s)$ is nondecreasing in $s$. By Lemma 4 we have that $\gamma_{s.a1l}\geq_{st}\gamma_{s.al}$. Thus, by Lemma 3 we have that $\big(E_{\gamma_{s.a1l}}E_{U}(g_{sl}(U)|A=a,S=s)-E_{\gamma_{s.a}}E_{U}(g_{sl}(U)|A=a,S=s)\big)\geq 0$.
N.2.4 Case 1b. LP-based bounds
Proposition 17.
Consider the setting of Figure 1(b). The LP-based bounds are
$\max_{s_{1},s_{2},l_{1},l_{2}\in\{0,1\}^{4}}\!\left\{1-\frac{1-p^{\,l}_{00.1l_{1}s_{1}}}{p^{\,l}_{10.0l_{2}s_{2}}}\right\}\leq VE(0)\leq\min_{s_{1},s_{2},l_{1},l_{2}\in\{0,1\}^{4}}\!\left\{1-\frac{p^{\,l}_{10.1l_{1}s_{1}}}{1-p^{\,l}_{00.0l_{2}s_{2}}}\right\},$
$\max_{s_{1},s_{2},l_{1},l_{2}\in\{0,1\}^{4}}\!\left\{1-\frac{1-p^{\,l}_{01.1l_{1}s_{1}}}{p^{\,l}_{11.0l_{2}s_{2}}}\right\}\leq VE(1)\leq\min_{s_{1},s_{2},l_{1},l_{2}\in\{0,1\}^{4}}\!\left\{1-\frac{p^{\,l}_{11.1l_{1}}}{1-p^{\,l}_{01.0l_{2}}}\right\},$
$\max_{s_{1},s_{2},l_{1},l_{2}\in\{0,1\}^{4}}\!\left\{1-\frac{1-p^{\,l}_{01.1l_{1}s_{1}}}{p^{\,l}_{10.0l_{2}s_{2}}}\right\}\leq VE_{T}\leq\min_{s_{1},s_{2},l_{1},l_{2}\in\{0,1\}^{4}}\!\left\{1-\frac{p^{\,l}_{11.1l_{1}s_{1}}}{1-p^{\,l}_{00.0l_{2}s_{2}}}\right\}.$
Proof. We again outline the main steps of the proof for deriving the LP-based bounds. Let $r_{y\mid a,l,s,b}$ denote the potential response of $Y^{a,m}$ given $A=a$, $S^{a,m=-1}=s$, $L=l$, and $B^{a,m=-1}=b$, for all $a,m\in\{0,1\}^{2}$ in the two-arm trial. As in Section 4.1, and for clarity, when conditioning on $A=a$ we write $B^{a,m=-1}$ simply as $B$ and $S^{a,m=-1}$ as $S$. Furthermore, let $q_{i.alsb}=\Pr(r_{y\mid a,l,s,b}=i)$ and define
$Q_{a,l,s,b}(a,m)=\{r_{y\mid a,l,s,b}:\,Y^{a,m}=1\ \text{given }A=a,L=l,S=s,B=b\}.$
Then, similarly to Section 4.1, and using the fact that $Y^{a,m}\perp_{\mathcal{T}_{VI}}(A,L,S)$, we can express the target estimand $E(Y^{a,m})$ as
$\begin{split}E(Y^{a,m})&=E(Y^{a,m}\mid A=a,L=l,S=s)\\ &=\pi^{\,l}_{0.als}E(Y^{a,m}\mid A=a,L=l,S=s,B=0)+\pi^{\,l}_{1.als}E(Y^{a,m}\mid A=a,L=l,S=s,B=1)\\ &=\pi^{\,l}_{0.als}\sum_{i\in Q_{a,l,s,0}}q^{\,l}_{i.als0}+\pi^{\,l}_{1.als}\sum_{i\in Q_{a,l,s,1}}q^{\,l}_{i.als1}.\end{split}$ (44)
The constraints and the remaining steps of the proof are identical to those described in Section 4.1, except that we condition, in addition to $S$, also on $L$.
N.2.5 Case 1b. Monotonicity-based bounds
Consider the following assumption.
Assumption 16.
($U,L,S$ monotonicity) Consider the setting of Figure 1(b). Suppose that $Y\perp_{\mathcal{T}_{VI}}M\mid A,B,L,S,U$ and one of the following conditions holds:
  1. 1.
    $\Pr(Y=1\mid A=a,B=1,L=l,S=s,U=u)$ and $\Pr(B=1\mid A=a,L=l,S=s,U=u)$ are both nondecreasing or both nonincreasing in $u$, for all $a$, $l$, and $s$.
  2. 2.
    One of $\Pr(Y=1\mid A=a,B=1,L=l,S=s,U=u)$ or $\Pr(B=1\mid A=a,L=l,S=s,U=u)$ is nondecreasing in $u$ and the other is nonincreasing in $u$, for all $a$, $l$, and $s$.
Proposition 18.
Consider the setting of Figure 1(b). Under Assumptions 8 and 16, the monotonicity-based bounds are
$\max_{l,s\in\{0,1\}^{2}}\!\left\{1-\frac{p_{1.1}}{p^{\,l}_{1.00ls}}\right\}$ $\;\leq\;$ $VE(0)$ $\;\leq\;\min_{l,s\in\{0,1\}^{2}}\!\left\{1-\frac{p^{\,l}_{1.10ls}}{p_{1.0}}\right\},$
$\max_{l,s\in\{0,1\}^{2}}\!\left\{1-\frac{p^{\,l}_{1.11ls}}{p_{1.0}}\right\}$ $\;\leq\;$ $VE(1)$ $\;\leq\;\min_{l,s\in\{0,1\}^{2}}\!\left\{1-\frac{p_{1.1}}{p^{\,l}_{1.01ls}}\right\},$
$\max_{l_{1},l_{2},s_{1},s_{2}\in\{0,1\}^{4}}\!\left\{1-\frac{p^{\,l}_{1.11l_{1}s_{1}}}{p^{\,l}_{1.00l_{2}s_{2}}}\right\}$ $\;\leq\;$ $VE_{T}$  $\;\leq\;1-\frac{p_{1.1}}{p_{1.0}}.$
Proof.
As a result from Appendix J, under Figure 1(b), $E\!\left(Y^{a,m=1}\right)$ can be written as
$\begin{split}E_{U}\big[E\!\left(Y\mid A=a,B=m,U=u,S=s\right)\big].\end{split}$ (45)
Since under Assumption 16, $Y\perp_{\mathcal{T}_{VI}}L\mid A,S,B,M,U$, we can also write that
$E(Y^{a,m=1})=E_{U}\big[E\!\left(Y\mid A=a,B=m,U=u,S=s,L=l\right)\big]$
Also, from Appendix J, $E(Y\mid A=a,B=1,S=s,L=l)$ can be written as
$\begin{split}E_{U}\bigg[E\!\left(Y\mid A=a,B=1,U=u,S=s,L=l\right)\frac{\Pr(B=1\mid A=a,U=u,S=s,L=l)}{\Pr(B=1\mid A=a,S=s,L=l)}\bigg].\end{split}$ (46)
Let $\nu(u)=\frac{\Pr(B=1\mid A=a,U=u,S=s,L=l)}{\Pr(B=1\mid A=a,S=s,L=l)}$. Then $\nu(u)$ is nondecreasing in $u$ since $\Pr(B=1\mid A=a,S=s,U=u,L=l)$ is nondecreasing in $u$, and since under Figure 1(b), $U\perp A,S,L$, we have
$E_{U}[\nu(U)]=E_{U|A,S,L}[\nu(U)]=1.$
Thus, by Lemma 1 we have
$\begin{split}E&\!\left(Y\mid A=a,B=1,S=s\right)-E\!\left(Y^{a,m=1}\right)\\ =&\operatorname{Cov}_{U}\!\left(E\!\left(Y\mid A=a,B=1,U,S=s,L=l\right),\nu(U)\right)\geq 0.\end{split}$ (47)
The rest of the proof is similar to the proof in Appendix J.
N.2.6 Case 1c. LP-based and monotonicity-based bounds
When $S$ is unobserved, as in Figure 1(c), we can apply Propositions 34 and their underlying assumptions by replacing $S$ with $L$, which yields the bounds stated in Proposition 16.
N.3 Case 2: when $L$ does not lead to narrower bounds
In this section, we present examples of covariates $L$ that do not satisfy the conditions of Proposition 14 and therefore cannot be used to obtain narrower bounds. Inspired by Cai et al. (2007), we classify the observed covariate $L$ into five scenarios, illustrated in Figure N.2.
Figures 2(a)2(b) depict scenarios in which $L$ is a confounder affecting both $B$ and $Y$. In Figure 2(b), $L$ is also influenced by $U$. An example corresponding to Figure 2(a) is age. In many clinical trials, including the ENSEMBLE2 trial, AEs are more common among younger participants (Hardt et al., 2022). As previously argued, an AE may affect participants’ beliefs. Moreover, age can influence exposure to the pathogen, as younger individuals tend to be more socially active. An example of $L$ that is a confounder and is also affected by $U$ is prior infection with the pathogen, which was measured, for instance, in the Pfizer COVID-19 vaccine trial (Polack et al., 2020). Tissot et al. (2021) found that prior COVID-19 infection is associated with increased risk of AEs (which may influence belief) and with reduced risk of reinfection.
Figure 2(c) illustrates a scenario in which $L$ is a prognostic factor for $Y$. Here, $L$ may represent baseline comorbidities that affect infection risk. For example, in the ENSEMBLE2 trial, comorbidities such as diabetes and high blood pressure were reported to influence the outcome (Hardt et al., 2022).
Figures 2(d)2(e) represent cases where $L$ is an intermediate variable between $B$ and $Y$. A simple example is when exposure $E$ is observed. In malaria vaccine trials, for instance, the use of bed nets was recorded; this behavior may be influenced by participants’ beliefs and, in turn, may affect infection risk (Chandramohan and others, 2021).
AALLBBUUYYMM
AALLBBUUYYMM
AALLBBUUYYMM
AALLBBUUYYMM
AALLBBUUYYMM
Figure N.2: DAGs illustrating five scenarios where the observed covariate $L$ cannot be used to obtain narrower bounds. For clarity, $S$ and $E$ are omitted.
Appendix O Additional results for the numerical example
@@
Figure O.1: The relationship between the magnitude of $\gamma_{B^{1}}$ and the width of the bounds for $\mathrm{VE}(0)$, $\mathrm{VE}(1)$, and $\mathrm{VE}_{T}$ under the setting of Figure 3(d). Assumption 81 is preserved when $\gamma_{B^{1}}>0$ and increasingly violated as $\gamma_{B^{1}}$ becomes negative, with $\gamma_{B^{0}}>0$ throughout. The parameters $\gamma_{B^{0}}$ and $\gamma_{B^{1}}$ control whether this assumption holds: when both were positive, the assumption is preserved, whereas more negative values of $\gamma_{B^{1}}$ correspond to stronger violations. Circles represent the true values. LP: bounds constructed via Proposition 5 and the results in Appendix E; Monotonicity: bounds constructed via Proposition 6. Triangles at the left boundary indicate bounds that extended below 0% but were truncated for display. No bounds are displayed when the estimated lower bound exceeds the upper bound, reflecting violation of the assumed monotonicity assumption.
@@
Figure O.2: The relationship between the magnitude of $\delta_{U}$ and the width of the bounds for $\mathrm{VE}(0)$, $\mathrm{VE}(1)$, and $\mathrm{VE}_{T}$ under the setting of Figure 3(c) ($\beta_{S}=0$). The case $\delta_{U}=0$ corresponded to the situation where our assumed DAG matched the true data-generating DAG. As $\delta_{U}$ increased, however, the true data moved further away from this assumed structure, since the effect of $U$ on $S$ became stronger (Figures 3(a) and 3(c)).Circles represent the true values. LP-based bounds constructed via Proposition 3 and the results in Appendix E. Triangles at the left boundary indicate bounds that extended below 0% but were truncated for display.
@@
Figure O.3: The relationship between the magnitude of $\delta_{U}$ and the width of the bounds for $\mathrm{VE}(0)$, $\mathrm{VE}(1)$, and $\mathrm{VE}_{T}$ under the setting of Figure 3(d). Here $\beta_{S}=log(2)$, so the data were generated with both $U$ and $S$ affecting $Y$, while the assumed DAG omitted these effects (Figure 3(a)). Circles represent the true values. LP-based bounds constructed via Proposition 3 and the results in Appendix E. Triangles at the left boundary indicate bounds that extended below 0% but were truncated for display. No bounds are displayed when the estimated lower bound exceeds the upper bound, reflecting violation of the assumed structure.
O.1 Bounds and corresponding bootstrap CIs
Table O.1: Empirical coverage of 95% percentile-bootstrap CIs for LP-based bounds
@@
Table O.2: Empirical coverage of 95% percentile-bootstrap CIs for monotonicity-based bounds
@@
Appendix P Additional results for the real data example
P.1 The full probabilities used to generate $B$ in the real data example
@@
P.2 Additional results for the real data example
Figure P.1: Estimated LP-based and monotonicity-based bounds under Figure 3(a). Each horizontal line shows the identified interval for $VE(0)$, $VE(1)$, and $VE_{T}$, stratified by method (LP vs. Monotonicity) and whether Assumption 8 1 was imposed (solid = no, dashed = yes). Triangles at the left boundary indicate bounds that extended below 0% but were truncated for display. No bounds are displayed when the estimated lower bound exceeds the upper bound, reflecting violation of the assumed structure.
Table P.1: Estimated VE bounds (%) under Figure 3(d). Cells show Estimate (95% CI). Values below 1000-1000 are reported as -$\infty$.